[figure: 9.7 cm]

THE VAN HOVE SCENARIO, P-WAVE PAIRING AND DISORDERED Sr_{2}RuO_{4}.

G. Litak^{a}, B.L. Györffy^{b}, J.F. Annett^{b}, K.I. Wysokiński^{c} aDepartment of Mechanics, Technical University of Lublin, Lublin, Poland bH.H. Wills Physics Laboratory, University of Bristol, Bristol, UK cInstitute of Physics, Maria Curie-Skłodowska University, Lublin, Poland

On the basis of a simple 2-dimensional lattice model we study the degradation of T_{C} with increasing disorder in p-wave superconductors. We deal with the electron-electron attraction in the Hartree-Fock-Gorkov Approximation and treat the disorder in the Coherent Potential Approximation. We highlight the influence of a van Hove singularity near E_{F} on the T_{C} vs. residual resistivity curve and comment on the relevance of our results to experiments on Sr_{2}RuO_{4}.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Grzegorz Litak

Address for correspondence : Department of Mechanics Technical University of Lublin Nadbystrzycka 36 20-618 Lublin, Poland

Email address : litak@tytan.umcs.lublin.pl